OpenStudy (anonymous):

when using the substitution method, how do you recognize an inconsistent system?

OpenStudy (whpalmer4):

You'll get a nonsense statement like \(0 = 1\) For example: \[x + y = 3\]\[x+y = 4\] From the first equation, \(y = 3-x\) substitute into second equation \[x+y=4\]\[x+(3-x) = 4\]\[3=4\]